A ball is thrown vertically upwards with an initial velocity u reaches maximum height in 5 sec.The ratio of distance travelled by the ball in the second and seventh second is? ( g=10m/s²)
1
Expert's answer
2020-11-05T07:33:03-0500

"V= V_0-gt; \text{for }t=5 \ V=0"

"V_0= gt= 50"

"S(t) = V_0t-\frac{gt^2}{2}"

"S(2)=50*2-\frac{10*2^2}{2}=80"

"S(7)=50*7 -\frac{10*7^2}{2}=" 105

"\frac{S(2)}{S(7)}=\frac{80}{105}=\frac{16}{21}"

Answer:"\ \frac{16}{21}"